OpenStudy (anonymous):

Consider the following equation: 4y 2−x 2−24y−4x + 16 = 0 Write this hyperbola in standard form.

OpenStudy (anonymous):

(2y-6)^2 -36 -(x+2)^2 +4 +16=0 (2y-6)^2 - (x+2)^2=16